A line contains the points (3, -1) and (-1, 2). Another line graphed in the same coordinate plane contains the points (2,0) and (-2,3).
Based on the slope of these lines, are they parallel, perpendicular or neither?

Found 2 solutions by rfer, solver91311:
Answer by rfer(12657)

m=(2--1)/(-1-3)
m=3/-4
m=-3/4
-----------------
m=(3-0)/(-2-2)
m=3/-4
m=-3/4
------------
parallel

Answer by solver91311(16877)



Use the slope formula on each pair of points:



where and are the coordinates of the given points.

If the slopes are equal, then the lines are parallel.

If the slopes are negative reciprocals, then the lines are perpendicular.

If the slopes are neither equal nor negative reciprocals, then the lines are neither parallel nor perpendicular.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism